OpenStudy (anonymous):

Ian's science book has 20 fewer pages than his math book. Ian's science book has 250 pages. Which equation can be used to calculate how many pages Ian's math book has? Let the variable p stand for the number of pages in the math book. A. 250 – 20 = p B. 250 = p + 20 C. 20 – p = 250 D. 250 = p – 20

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (sagewilson):

It's D, not A "Ian's science book has 20 fewer pages than his math book." making his math book 270 pages. 250 – 20 = p is out of the question. Knowing that p equals 270, 250 = p + 20, wouldn't make any sense. 20 – p = 250 would leave you with a negative number, not 250. So D is the correct answer.